# GOOD: Towards Domain Generalized Orientated Object Detection

### v1

@@ 4.3 Ablation Studies
On Different Distance Metric Table 8 investigates the impact of different distance metric on $\mathcal{L}_{SC}$ (in Eq. 10). The JSD distance metric in $\mathcal{L}_{SEC}$ is compared with four commonly-used distance metrics, namely, $l$-2 distance and K-L distance. The style consistency learning from $l$-2 and K-L distance is less significant than the JSD distance.
@@
4.4 Visualization
Fig. 6 visualizes some detection results on unseen target domains from the ReDet baseline and the proposed GOOD. The proposed GOOD shows more reasonable and more precise rotated bounding boxes on these unseen target domains. More visualization results in the supplementary materials.
@@ 5 Conclusion
In this paper, we make an early exploration to learn domain generalized oriented object detection from a single source domain. It is particularly challenging as the detector need to represent not only stable content but also precise orientation under the cross-domain style variation. In the proposed GOOD, the CLIP-driven style hallucination is constrainted by the proposed rotation-aware content consistency learning and style-consistency learning, which allows the model to learn generalizable rotation and content representation.
Extensive experiments on a series of domain generalization settings show that the proposed GOOD has significantly stronger generalization than existing state-of-the-art oriented object detectors.
Limitation & Future Work. We focus on existing domain generalized detection settings, where the cross-domain object categories are consistent. In the future, we will GOOD under open-world scenarios, which can further discover unknown categories from different domains.

### v2

@@ 4.5.5 On Different Distance Metric
[…]
To this end, Table 13 investigates the impact of different distance metric on $\mathcal{L}_{SEC}$ (in Eq. 10). The JSD distance metric in $\mathcal{L}_{SEC}$ is compared with two commonly-used distance metrics, namely, $l$-2 distance and K-L distance. It leads to an mAP improvement of 1.06% and 0.65% on the SODA and FAIR1M unseen target domains than the $l$-2 distance metric. Similarly, it leads to an mAP improvement of 0.67% and 0.15% on both unseen target domains than the K-L distance metric.
To conclude, the style consistency learning from $l$-2 and K-L distance is less significant than the JSD distance. As discussed in Sec. 3.4, JSD distance is more capable to handle scenarios when the category scores of pre- and post- hallucinated samples vary greatly in terms of the numerical value, thus yielding better generalization performance.
[…]
@@
4.6 Visualization
To study if the proposed GOOD is able to visually predict more reasonable and more precise rotated bounding boxes on multiple object categories from these unseen target domains, we display some prediction results on large-scale aerial images. Both the results from the ReDet Han et al. (2021b) baseline and the proposed GOOD are displayed, which we denote as Baseline and Ours, respectively.
@@ 4.6 Visualization
The first two columns in Fig. 8 provide the visual results on large-scale aerial images. In the first scene, compared with the baseline, the proposed GOOD predicts more precise orientation for baseball field. In the second scene, compared with the baseline, the proposed GOOD predicts more precise orientation for tennis court. Besides, in both scenes, more large vehicle objects are detected by GOOD.
[…]
The second two columns in Fig. 8 show some visual predictions when the baseline and the proposed GOOD is generalized to the FAIR1M and SODA target domain, respectively. The proposed GOOD not only shows more precise orientation predictions on airplanes, but also more capable to discern the vehicle category.
[…]
Fig. 9 provides the visual results on this cross-domain setting. In the third road scene, compared with the baseline, the proposed GOOD significantly reduces the false alarms of small vehicles (upper left of the image) and large vehicles (at the bottom of the image). In the fourth urban scene, compared with the baseline, the proposed GOOD reduces some false alarms of small vehicles along the road and large vehicles at the center of the image.
@@ 5 Conclusion
In this paper, we make an early exploration to learn domain generalized oriented object detection from a single source domain. It is particularly challenging as the detector need to represent not only stable content but also precise orientation under the cross-domain style variation. In the proposed GOOD, the CLIP-driven style hallucination is constrained by the proposed rotation-aware content consistency learning and style-consistency learning, which allows the model to learn generalizable rotation and content representation. Experiments on a series of domain generalization settings show that the proposed GOOD has significantly stronger generalization than existing state-of-the-art oriented object detectors.
Limitation & Future Work.
One may argue the proposed GOOD only follows the two-stage paradigm. But compared with the one-stage paradigm, it provides a more precise description between orientation and region of interests. It also provides more access to enrich the style and constrain the content consistency from shallow to deep. On the other hand, we focus on existing domain generalized detection settings, where the cross-domain object categories are consistent. In the future, we will advance GOOD to open-world scenarios, which can further discover unknown categories from different domains.
Acknowledgment
This work was supported by the National Natural Science Foundation of China under contracts No.U22B2011 and No.62325111.